# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v3

@@ 6 Deep model with the spectral truncation kernels
As we discussed in Section 4, there is a tradeoff between the representation power and the model complexity. In fact, there are two directions of the representation power: one is about the dependency of the outputs on the inputs discussed in Section 4, and the other is about the interactions along the data function domain. One shortcoming of the spectral truncation kernel is that to obtain a model with a high representation power in the sense of the dependency of the outputs on the inputs, the truncation parameter $n$ have to be large, which results in making the interactions along the data function domain weak. To gain the two types of representation powers at the same time, we consider the following deep model $f$ with $L$ spectral truncation kernels $k_{n}^{1},\ldots,k_{n}^{L}$ and learnable parameters $c_{i}^{1},\ldots,c_{i}^{L}\!\!\in\!\mathcal{A}$:
$f(x)=\prod_{j=1}^{L}\bigg{(}\sum_{i=1}^{N}k_{n}^{j}(x,x_{i})c^{j}_{i}\bigg{)}.$ (6)
We consider the case where $c_{i}^{j}$ in Eq. (6) is represented as a finite sum of functions in $\mathcal{A}$. In the following case, the representation power of this model grows exponentially.
@@
The exponential growth of the representation power is also observed for neural networks (Hanin & Rolnick, 2019). Deep learning with kernels has also been proposed (Laforgue et al., 2019; Hasimoto et al., 2023). Whereas the growth is obtained by the composition for these existing frameworks, it is obtained by the product for our case. Unlike the case of the existing frameworks, the representation power regarding $\mathcal{A}$ does not become high by the composition of functions.
@@
The function $f$ is in the RKHM associated with the kernel $\prod_{j=1}^{L}k^{j}$. Since the Rademacher complexity bound depends on the norm of the kernel, it can also grow exponentially with respect to $L$. Thus, for the deep setting, in the same manner as Theorem 16, we still have the tradeoff between the representation power and the model complexity. In this case, it is controlled by both $n$ and $L$. The exponential growth of the Rademacher complexity bound (thus, the generalization bound) with respect to the number of layers is also observed for neural networks (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018). Investigating how we can take advantage of the deep setting in the sense of the generalization is future work.
@@ 7.1 Experiment with synthetic data
We first observed the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, and ${k}_{n}^{\operatorname{sep},q}$ with different values of $n$ and compared them to the commutative kernels $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ (corresponding to $n=\infty$). We considered a regression task with synthetic data. For $i=1,\ldots,N$, we generated input training samples as $x^{i}(z)=[\sin(0.01iz)+0.01\xi_{z,i},\cos(0.01iz)+0.01\eta_{z,i}]$. We set the target function $f$ as $f(x)(z)=3(\sin(\cos(\int_{z-\Delta}^{z+\Delta}x_{1}(t)\mathrm{d}t+\int_{z-\Delta}^{z+\Delta}x_{2}(t)\mathrm{d}t)))$, where $\Delta=2\pi/30$, and generated training output samples $y^{i}(z)=f(x)(z)+0.001\xi_{z,i}$. Here, $\xi_{z,i}$ and $\eta_{z,i}$ were sampled independently from the Gaussian distribution with mean 0 and standard deviation 1. We set the sample size $N$ as 1000. For the kernel ${k}_{n}^{\operatorname{poly},q}$, we set $q=1$ and $\alpha_{1}=1$. Note that ${k}^{\operatorname{poly},q}$ is the linear kernel in this case. For the kernel $\hat{k}_{n}^{\operatorname{prod},q}$, we set $q=1$, $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=2\pi\mathrm{e}^{-|x-y|^{2}}$. We set $\beta_{n}=1$ for $n<\infty$ and $\beta_{\infty}=0$. For the kernel $k_{n}^{\operatorname{sep},q}$, we set $\tilde{k}(x,y)=\mathrm{e}^{-0.1\cdot 2\pi^{2}(\|x_{1}-y_{1}\|_{L^{2}(\mathbb{T})}^{2}+\|x_{2}-y_{2}\|_{L^{2}(\mathbb{T})}^{2})}$, $q=2$, and $a_{1}(z)=a_{2}(z)=2\pi\mathrm{e}^{\sin z}$. We estimated $f$ using kernel ridge regression. We applied the cross-validation grid search to find an optimal regularization parameter $\lambda$. Figure 2 (a-c) illustrates the test error. We can see that for ${k}_{n}^{\operatorname{poly},q}$, the test error is the smallest when $n=32$. It becomes large if $n$ is larger or smaller than $32$. This is because when $n=32$, the empirical error and the complexity in Eq. 3 are balanced, and the best possible expected error is obtained. We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the typical commutative kernel ($n=\infty$). We have similar results for the other two kernels. Note that $\hat{k}_{\infty}^{\operatorname{prod},q}(x,y)(z)=\hat{k}^{\operatorname{prod},q}(x,y)(z)$ only depends on $x(z)$ and $y(z)$. However, $f(x)(z)$ depends on $x(t)$ for $t\in[z-\Delta,z+\Delta]$. Thus, the test error becomes large when $n=\infty$. In addition, as discussed in Remark 15, $\hat{k}_{n}^{\operatorname{prod},q}$ may not be positive definite. However, although we set $\beta_{n}$ as a small value, all the eigenvalues of the Gram matrix are positive in this case. See Appendix E.1 for more details.
@@
We also compared the performance and the computational time of RKHMs associated with the proposed kernels to a vvRKHS with the same setting as above. For the vvRKHS, we used the nonseparable kernel $k(x,y)=\tilde{k}(x,y)[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$, where we replaced $S_{n}(\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j}))$ with the nonseparable matrix $[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$. This nonseparable kernel is proposed by Lim et al. (2015). Table 1 shows the result. We can see that the RKHMs outperform the vvRKHS, and the computational time for the RKHMs are smaller than that for the vvRKHS. Note that we have already confirmed the proposed kernels outperform the existing separable kernel $k_{\infty}^{\operatorname{sep},q}$ in the above experiment (See Figure 2 (c)).
@@
Finally, we observed the behavior of the deep model. We generated data in the same manner as above, but we set $N=300$. We used the kernel ${k}_{n}^{\operatorname{prod},q}$ and set $L=1,2,3$, $\tau_{1}=\sqrt{2}$, $\tau_{2}=\sqrt{3}$, and $\tau_{3}=\sqrt{5}/2$. We set $n$ so that the numbers of parameters are the same for all values of $L$. We used the loss function $\|\sum_{i=1}^{N}|f(x_{i})-y_{i}|_{k}\|_{L^{2}(\mathbb{T})}+0.1\||f|_{k}\|_{L^{2}(\mathbb{T})}$ for the deep model $f$ defined in Eq. (6). Figure 2 (d) shows the result. We can see that as the number of layers increases, we can obtain a higher performance model even with the same number of parameters.
@@
(c) $k=k_{n}^{\operatorname{sep},q}$
(d) Deep approach
$n=30$ $n=40$ $n=50$ $n=60$ $n=70$ $n=80$ $n=\infty$
(e) Output images
Figure 2: (a-c) Test error of the regression task with different values of $n$. (Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (d) Test error of the regression task with deep approach with different $n$ and $L$. The parameters $n$ and $L$ are chosen so that the numbers of parameters are the same for all the cases. (e) Output images of the image recovering task with different values of $n$.
@@
RKHM, $k=k_{n}^{\operatorname{sep},q}$ ($n=10$) 0.0385$\pm$5.04388×$\times 10^{-5}$ 22.79$\pm$ 0.5742
vvRKHS 0.0774$\pm$7.86 $10^{-5}$ 570.4$\pm$14.87
@@ 7.2 Experiment with MNIST
We consider an image recovering task with MNIST (LeCun et al., 1998). For $i=1,\ldots,N$, we generated input training samples by setting all the $8\times 8$ pixels in the middle of the images as 00. We set the output samples as the original images. We flatten the 2-dimensional $28\times 28$ image to a vector in $\mathbb{R}^{784}$ and regard it as a discretized function on the space of pixels. We tried to estimate a function that transforms an image with the missing part into the original image. We applied kernel ridge regression with the regularization parameter $\lambda=0.01$. We set the sample size $N$ as $200$, and used the kernel $k_{n}^{\operatorname{prod},q}$ with $q=1$ and $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=2\pi\mathrm{e}^{-0.1|x-y|^{2}}$. We set $\beta_{n}=0.01$ for $n<\infty$ and $\beta_{\infty}=0$. Figure 2 (e) shows the output images with different values of $n$. When $n=\infty$, i.e., the commutative kernel, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image. When $n=70$, we can recover the image the clearest.
8 Conclusion and Limitation

### v4

@@ 6 Deep model with the spectral truncation kernels
As we discussed in Subsection 3.1, the truncation parameter $n$ controls the global and local dependencies of the output on the input. To make the framework more flexible and capture both global and local dependencies at the same time, we consider the following deep model $f$ with $L$ spectral truncation kernels $k_{n_{1}}^{1},\ldots,k_{n_{L}}^{L}$ and learnable parameters $c_{i}^{1},\ldots,c_{i}^{L}\!\!\in\!\mathcal{A}$:
$f(x)=\prod_{j=1}^{L}\bigg{(}\sum_{i=1}^{N}k_{n_{j}}^{j}(x,x_{i})c^{j}_{i}\bigg{)}.$ (8)
If we set a different value of $n_{j}$ for each layer $j$, then we can capture different level of global and local dependencies for each layer, which enables us to capture both global and local dependencies.
To learn the deep model, we need to represent $c_{i}^{j}$ in Eq. (8) as a finite sum of functions in $\mathcal{A}$. The following proposition provides an appropriate way to parameterize $c_{i}^{j}$ in the sense that the representation power of the model grows exponentially. For simplicity, we consider the case where $n_{j}=n$ for $j=1,\ldots,L$ in the following proposition. However, we can derive the same result for the case of $n_{j}\neq n_{l}$ for $j\neq l$.
@@
The exponential growth of the representation power is also observed for neural networks (Hanin and Rolnick, 2019). Deep learning with kernels has also been proposed (Laforgue et al., 2019; Hasimoto et al., 2023). Whereas the growth is obtained by the composition for these existing frameworks, it is obtained by the product for our case. Unlike the case of the existing frameworks, the representation power regarding $\mathcal{A}$ does not become high by the composition of functions.
@@
The function $f$ is in the RKHM associated with the kernel $\prod_{j=1}^{L}k_{n_{j}}^{j}$. Since the Rademacher complexity bound depends on the norm of the kernel, it can also grow exponentially with respect to $L$. Thus, for the deep setting, in the same manner as Theorem 17, we still have the tradeoff between the representation power and the model complexity. In this case, it is controlled by both $n$ and $L$. The exponential growth of the Rademacher complexity bound (thus, the generalization bound) with respect to the number of layers is also observed for neural networks (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018). Investigating how we can take advantage of the deep setting in the sense of the generalization is future work.
7 Applications: Extracting Local and Global Dependencies of functions
There are many potential applications of the spectral truncation kernels. We list two examples here.
Time-series data analysis
We can regard a time-series as a function on a time space. In many cases, a state at a certain time $z$ is influenced strongly by another state at the same time $z$, but also by the state around the time $z$. Since commutative kernels focus only on local information, we cannot describe these two states with commutative kernels. On the other hand, since separable kernels focus only on global information, we cannot describe the relationship of these two states at each time $z$. By applying the proposed kernels, we can extract global information, but also can focus on local information.
Operator learning
In the framework of operator learning, we obtain a solution of a partial differential equation as an output from an input function (such as initial condition or parameter of the equation). Thus, we construct a model where both of the input and output are functions. Applying kernel methods to operator learning has been proposed (Batlle et al., 2024). We can construct the model by solving a kernel ridge regression task. Another well-known operator learning method is neural operator. In the framework of neural operators, we apply integral operators to extract global information and apply local linear operators and local activation functions to extract local information. The proposed kernel enables us to do similar procedures for the operator learning with kernels. By considering the product of multiple proposed kernels with different values of $n$ or deep model with the proposed kernels with different values of $n$, we can extract both global and local information; we can extract global information using the kernel with small $n$ and extract local information using the one with large $n$ in the model. We will numerically confirm the applicability of the spectral truncation kernels to operator learning in Subsection 8.3.
@@ 8.1 Experiment with synthetic data
We first observed the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$ and $\hat{k}_{n}^{\operatorname{prod},q}$ with different values of $n$ to observe that the truncation parameter $n$ controls global and local dependencies of the output on the input, as discussed in Subsection 3.1. We considered a regression task with synthetic data. For $i=1,\ldots,N$, we generated input training samples as $x^{i}(z)=[\sin(0.01iz)+0.01\xi_{z,i},\cos(0.01iz)+0.01\eta_{z,i}]$. We set the target function $f$ as $f(x)(z)=3(\sin(\cos(\int_{z-\Delta}^{z+\Delta}x_{1}(t)\mathrm{d}t+\int_{z-\Delta}^{z+\Delta}x_{2}(t)\mathrm{d}t)))$, where $\Delta=2\pi/30$, and generated training output samples $y^{i}(z)=f(x)(z)+0.001\xi_{z,i}$. Here, $\xi_{z,i}$ and $\eta_{z,i}$ were sampled independently from the Gaussian distribution with mean 0 and standard deviation 1. We set the sample size $N$ as 1000. For the kernel ${k}_{n}^{\operatorname{poly},q}$, we set $q=1$ and $\alpha_{1}=1$. Note that ${k}^{\operatorname{poly},q}$ is the linear kernel in this case. For the kernel $\hat{k}_{n}^{\operatorname{prod},q}$, we set $q=1$, $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=2\pi\mathrm{e}^{-|x-y|^{2}}$. We set $\beta_{n}=1$ for $n<\infty$ and $\beta_{\infty}=0$. We estimated $f$ using kernel ridge regression with the regularization parameter $\lambda=0.01$. We generated 1000 input test samples in the same manner as the input training samples, and evaluated the test error $(1/N)\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimation of $f$ and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively. For $x^{i}$, $y^{i}$, and $\tilde{x}^{i}$, we discretized each function with 30 equally spaced points on $\mathbb{T}$. We applied the cross-validation grid search to find an optimal regularization parameter $\lambda$. Figure 3 (a,b) illustrates the test error. We can see that for ${k}_{n}^{\operatorname{poly},q}$, the test error is the smallest when $n=32$. It becomes large if $n$ is larger or smaller than $32$. This is because when $n=32$, the information of the global and that of the local dependencies are balanced, and the best possible error is obtained. We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the typical commutative kernel ($n=\infty$). We have similar results for ${k}_{n}^{\operatorname{prod},q}$. Note that $\hat{k}_{\infty}^{\operatorname{prod},q}(x,y)(z)=\hat{k}^{\operatorname{prod},q}(x,y)(z)$ only depends on $x(z)$ and $y(z)$. However, $f(x)(z)$ depends on $x(t)$ for $t\in[z-\Delta,z+\Delta]$. Thus, the test error becomes large when $n=\infty$. In addition, as discussed in Remark 16, $\hat{k}_{n}^{\operatorname{prod},q}$ may not be positive definite. However, although we set $\beta_{n}$ as a small value, all the eigenvalues of the Gram matrix are positive in this case. To investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0. Figure 3 (c) shows the result. The index $i$ is determined in the descendant order with respect to the value of $\lambda_{i}$. Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
@@
We also compared the performance and computational time with the proposed kernels to the existing operator-valued kernel with the same setting as above to show the advantage of the proposed kernel in the sense of the computational efficiency and performance. For the existing kernel, we used the nonseparable kernel $k(x,y)=\tilde{k}(x,y)[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$, where we replaced $S_{n}(\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j}))$ with the nonseparable matrix $[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$. This kernel is the combination of the separable and transformable kernels and proposed by Lim et al. (2015). Table 2 shows the result. We can see that the proposed kernels outperform the existing typical nonseparable kernel, and the computational time with the proposed kernel is smaller than that with the existing kernel.
@@
Finally, we observed the behavior of the deep model to show its high representation power. We generated data in the same manner as above, but we set $N=300$. We used the kernel ${k}_{n}^{\operatorname{prod},q}$ and set $L=1,2,3$, $\tau_{1}=\sqrt{2}$, $\tau_{2}=\sqrt{3}$, and $\tau_{3}=\sqrt{5}/2$. We set $n_{j}=n$ for $j=1,\ldots,L$ for simplicity and set $n$ so that the numbers of parameters are the same for all values of $L$. We used the loss function $\|\sum_{i=1}^{N}|f(x_{i})-y_{i}|_{k}\|_{L^{2}(\mathbb{T})}+0.1\||f|_{k}\|_{L^{2}(\mathbb{T})}$ for the deep model $f$ defined in Eq. (8). Figure 4 shows the result. We can see that as the number of layers increases, we can obtain a higher performance model even with the same number of parameters.
@@
(c) Eigenvalues
Figure 3: (a,b) Test error of the regression task with different values of $n$. (Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (c) Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the regression task. (Average value of results of five different runs. The error bar represents the standard deviation.)
@@
vvRKHS (combination of transformable
and separable kernels)
0.0774$\pm$7.86 $10^{-5}$ 570.4$\pm$14.87
Figure 4: Test error of the regression task with deep approach with different $n$ and $L$. The parameters $n$ and $L$ are chosen so that the numbers of parameters are the same for all the cases. (Average value $\pm$ the standard deviation of three independent runs.)
$n=30$ $n=40$ $n=50$ $n=60$ $n=70$ $n=80$ $n=\infty$
Figure 6: Output images of the image recovering task with different values of $n$.
@@ 8.2 Experiment with MNIST
To check the performance of the proposed kernels for practical applications, we consider an image recovering task with MNIST (LeCun et al., 1998). For $i=1,\ldots,N$, we generated input training samples by setting all the $8\times 8$ pixels in the middle of the images as 00. We set the output samples as the original images. We flatten the 2-dimensional $28\times 28$ image to a vector in $\mathbb{R}^{784}$ and regard it as a discretized function on the space of pixels. We tried to estimate a function that transforms an image with the missing part into the original image. We applied kernel ridge regression with the regularization parameter $\lambda=0.01$. We set the sample size $N$ as $200$, and used the kernel $k_{n}^{\operatorname{prod},q}$ with $q=1$ and $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=2\pi\mathrm{e}^{-0.1|x-y|^{2}}$. We set $\beta_{n}=0.01$ for $n<\infty$ and $\beta_{\infty}=0$. We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $(1/N)\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively. Figure 5 (a) shows the test error. We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$. In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0. Figure 5 (b) shows the result. The index $i$ is determined in the descendant order with respect to the value of $\lambda_{i}$. Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case. Furthermore, Figure 6 shows the output images with different values of $n$. When $n=\infty$, i.e., the commutative kernel, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image. When $n=70$, we can recover the image the clearest.
8.3 Experiment for operator learning
To show the availability of the proposed kernels to operator learning, we conducted an experiment. Consider the following Burgers’ equation on $(0,1]\times(0,1]$ for $w(x,t)$.
$\frac{\partial w}{\partial t}+w\frac{\partial w}{\partial x}$ $=\nu\frac{\partial^{2}w}{\partial x^{2}}\qquad x\in(0,1)\times(0,1],$
$w(x,0)$ $=u(x)\qquad x\in(0,1),$
where $\nu=0.1$. We learned the operator mapping the initial condition $u$ to $w(\cdot,1)$, the solution at time $t=1$, by kernel ridge regression explained in Section 5. The training data and test data are generated by sampling the initial condition $u$ from the Gaussian distribution with mean 0 and the covariance $625(-\Delta+25I)^{-2}$. We used equally spaced $128$ points for computing the integral related to the input functions. The number of training samples is 1000 and that of test samples is 200. The same problem is also considered by Batlle et al. (2024). We used the same $\hat{k}_{n}^{\operatorname{prod},q}$ kernel as Subsection 8.1 with $n=5$. We set the regularization parameter $\lambda$ as $0.01$.
Figure 7 shows the true solution $w(\cdot,1)$ and the predicted function $v$ by using the proposed kernel for the same test input function $u$ (initial condition). Figure 8 shows the error $v-w(\cdot,1)$ of these two functions. The mean value of the pointwise error $(1/128)\sum_{i=1}^{128}|v(x_{i})-w(x_{i},1)|$ over all the test samples is $0.00157\pm 7.079\times 1e^{-6}$ (average $\pm$ standard deviation over five independent runs). We can see that the proposed kernel combined with kernel ridge regression properly predict the solution of the problem, and compared to the results in Subsection 4.3.1 by Batlle et al. (2024), the proposed kernel outperforms these existing results.
Figure 7: The true solution $w(\cdot,1)$ and the predicted function $v$ of the Burgers’ equation.
Figure 8: Pointwise error $v-w(\cdot,1)$ of the true solution $w(\cdot,1)$ and the predicted function $v$ of the Burgers’ equation.
9 Generalization of the Spectral Truncation Kernels
The proposed spectral truncation kernels have potential possibilities of constructing kernels for more general situations.
9.1 Generalization to other $C^{*}$-algebras
Continuous functions on high-dimensional torus (Leimbach and van Suijlekom, 2024)
[…]
Continuous functions on the sphere (Rieffel, 2004)
9.2 Generalization using other integral kernels
We can generalize the spectral truncation kernel to a kernel on a measure space $\mathcal{X}$ by replacing the function $F_{n}^{2q,P}$ in Eq. (4) with another function $G:\Theta\times\mathcal{X}^{2q}$ that satisfies $G(\theta,z)\geq 0$ for any $\theta\in\Theta$ and $z\in\mathcal{X}^{2q}$. Here, $\Theta$ is the space of parameters. For example, we can set $\Theta=\mathbb{R}_{+}$, the set of positive real values, $\mathcal{X}=\mathbb{R}^{d}$, and $G(\sigma,z)=\mathrm{e}^{-\sigma\|z\|^{2}}$. Then, similar to the spectral truncation kernel, the convolution with $G(\sigma,z-\cdot)$ captures global information if $\sigma$ is small and captures local information if $\sigma$ is large. In this case, since the parameter space $\Theta$ is continuous, we can use an optimization method to learn a best possible parameter $\sigma$ from data.
10 Conclusion and Discussion